find the molecular weight of a gass
mass = 22.0 g
pressure = 856 torr
temp = 25.0°C
volume of container = 3.00 L
Answer #1

mass of gas [m]= 22.0 g

pressure [P] = 856 torr =[856/760] atm = 1.126 atm

temperature [T] = 25.0 +273.0 = 298 K

volume [V] = 3.00 L

gas constant R=0.0821 L.atm.K-1.mol-1

we know the ideal gas equation as PV = nRT where n=mass/molecular weight

substituting the values in ideal gas equation, we get

1.162*3.00 = [22g / molecular weight]*0.0821*298

or 3.486 = [22g / molecular weight] * 24.4658

or 3.486/24.4658 = 22g / molecular weight

or 0.1425 = 22g / molecular weight

or molecular weight = 22 g / 0.1425 = 154.386 g.mol-1